Session: June 2008 (session 2)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Shanghai (Antao at Sat May 31 05:33:38 2008) Success :
Onsala (Michael Lindqvist at Sat May 31 06:39:25 2008) Minor failures :
Problem with "noisy" RCP channel back (see em059c). Have not been able to find any problems with the receiver (yet).
Medicina (Morsiani at Sat May 31 06:40:10 2008) Success :
Effelsberg (Uwe Bach at Mon Jun 2 08:35:05 2008) Success :
Noto (P. Cassaro at Tue Jun 10 09:42:01 2008) Success :
Jodrell_Bank (Alastair Gunn at Tue Jun 10 23:43:06 2008) Success :
The MkII telescope was used (Jb2).
No known problems.
Cambridge (Alastair Gunn at Tue Jun 10 23:43:54 2008) Minor failures : RECO
Disk pack NAIC-004 stopped recording. Rebooted Mk5 and changed to pack WSRT-019 from about 03:15 UT. No other known problems.